b) For 25 army personnels, line of regression of weight of kidneys (Y) on weight of 

heart (X ) is Y = .0 399X + .6 934 and the line of regression of weight of heart on 

weight of kidney is X − .1 212Y + .2 461= .0 Find the correlation coefficient between 

X and Y and their mean values.


1
Expert's answer
2021-04-15T06:50:21-0400

Rewrite the regression equations uniformly


"Y = 0.0 399X + 0.6 934"




"X = 0 .1 212Y - 0.2 461"


The basic formulas are


"Y = r\dfrac{s_y}{s_x}X -r\dfrac{s_y}{s_x}\bar{x}+\bar{y}"

"X = r\dfrac{s_x}{s_y}Y -r\dfrac{s_x}{s_y}\bar{y}+\bar{x}"

Then


"r\dfrac{s_y}{s_x}=0.0399"

"r\dfrac{s_x}{s_y}=0.1212"

Hence


"r^2\dfrac{s_y}{s_x}\cdot\dfrac{s_y}{s_x}=0.0399(0.1212)"

"r^2=0.00483588"

"r=\sqrt{0.00483588}\approx0.06954049"

"\begin{cases}
 -0.0399\bar{x}+\bar{y}=0.6934 \\
 -0.1212\bar{y}+\bar{x}=-0.2461
\end{cases}"

"\begin{cases}
 \bar{y}=0.6934+0.0399\bar{x} \\
 -0.1212(0.6934+0.0399\bar{x} )+\bar{x}=-0.2461
\end{cases}"

"\begin{cases}
 \bar{y}=0.6934+0.0399\bar{x} \\
 \bar{x}=-\dfrac{0.16205992}{0.99516412}
\end{cases}"

"\begin{cases}
 \bar{y}=0.6869 \\
 \bar{x}=-0.1628
\end{cases}"

"r=0.0695"